# Mirror Descent Algorithms with Nearly Dimension-Independent Rates for Differentially-Private Stochastic Saddle-Point Problems

### v1

@@ Abstract
We study the problem of differentially-private (DP) stochastic (convex-concave) saddle-points in the polyhedral setting. We propose $(\varepsilon,\delta)$-DP algorithms based on stochastic mirror descent that attain nearly dimension-independent convergence rates for the expected duality gap, a type of guarantee that was known before only for bilinear objectives. For convex-concave and first-order-smooth stochastic objectives, our algorithms attain a rate of $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{1/3}$, where $d$ is the dimension of the problem and $n$ the dataset size. Under an additional second-order-smoothness assumption, we improve the rate on the expected gap to $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{2/5}$. Under this additional assumption, we also show, by using bias-reduced gradient estimators, that the duality gap is bounded by $\log(d)/\sqrt{n}+\log(d)/[n\varepsilon]^{1/2}$ with constant success probability. This result provides evidence of the near-optimality of the approach. Finally, we show that combining our methods with acceleration techniques from online learning leads to the first algorithm for DP Stochastic Convex Optimization in the polyhedral setting that is not based on Frank-Wolfe methods. For convex and first-order-smooth stochastic objectives, our algorithms attain an excess risk of $\sqrt{\log(d)/n}+\log(d)^{7/10}/[n\varepsilon]^{2/5}$, and when additionally assuming second-order-smoothness, we improve the rate to $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$. Instrumental to all of these results are various extensions of the classical Maurey Sparsification Lemma, which may be of independent interest.
@@ 1 Introduction
Protecting privacy of users’ data has become a key desideratum in machine learning. Such a requirement is fundamental in situations involving e.g. medical or financial data. In this context, differential privacy (DP) has become the de facto standard for private data analysis, and it has been intensively studied for nearly 2 decades (Dwork et al., 2006b, 2014).
Recently, an emerging line of work has focused on solving stochastic saddle-point [^1] We will restrict ourselves to the convex-concave setting, and therefore omit this from the problem denomination (SSP) problems under DP (Boob and Guzmán, 2023; Zhang et al., 2022a; Yang et al., 2022; Kang et al., 2022; Bassily et al., 2023). While this problem has been primarily motivated by group fairnes in machine learning (Jagielski et al., 2019; Li et al., 2019; Lowy et al., 2023), its generality may be of use in problems as varied as learning under distributional robustness (Mohajerin Esfahani and Kuhn, 2018) and synthetic data generation (Gaboardi et al., 2014). In relation to the latter problem, a long list of works has explored the idea of using zero-sum games (corresponding to bilinear saddle-points with polyhedral feasible sets) for private query answering and private synthetic data generation (e.g.,  Hardt and Rothblum (2010); Hardt et al. (2012); Hsu et al. (2013); Gaboardi et al. (2014)).
@@
Under privacy constraints, there is a major discrepancy in the literature between the bilinear and convex-concave settings. While bilinear problems have convergences rates –in terms of the expected duality gap– which scale only poly-logarithmically with the ambient dimensions (e.g., Hardt and Rothblum (2010)), all known rates for general convex-concave objectives scale polynomially with the dimension (e.g., Bassily et al. (2023)). This can be attributed to the fact that all existing results for convex-concave settings work under Euclidean geometry, which assumes $\ell_{2}$ bounded feasible sets and objectives that are Lipschitz and smooth with respect to $\|\cdot\|_{2}$. Therefore, known polynomial-in-the-dimension lower bounds for DP Stochastic Convex Optimization (SCO) in the Euclidean setting (Bassily et al., 2014) apply. Recently, (Asi et al., 2021b; Bassily et al., 2021a) showed that it is possible to avoid the lower bounds for DP-SCO under additional geometric structure on the problem, which we refer to as the polyhedral setting. It consists of privately minimizing an stochastic objective which is Lipschitz and smooth with respect to $\|\cdot\|_{1}$ over a polyhedral feasible set. Inspired by these recent developments, in this work we extend the scope of polyhedral settings to the class of convex-concave, Lipschitz and smooth [^2] We only consider smooth objectives because even for DP-SCO in the polyhedral setting, smoothness is a necessary condition to get nearly dimension-independent rates (Asi et al., 2021b). SSP problems (see Section 4 for a formal problem statement). Below we summarize our results and main contributions.
Summary of Contributions.
We prove the first nearly-dimension independent convegence rates for convex-concave DP-SSP in the polyhedral setting. Furthemore, for (nearly) quadratic objectives, our rates are optimal. In addition, our mirror-descent based algorithms achieve nearly-dimension independent convegence rates for DP-SCO in the polyhedral setting. Moreover, our work is the first that does not use a Frank-Wolfe based method to get this type of guarantee. Table 1 summarizes the obtained rates.
@@
  • We propose two algorithms for DP-SSP, based on (entropic) Stochastic Mirror Descent (SMD). As opposed to previous results for bilinear settings where privatization by sampling from vertices leads to unbiased gradient estimators (Hsu et al., 2013), the main challenge we encounter is controlling the bias in the more general (stochastic) convex-concave setting. The first algorithm (Section 4.1) controls bias by simply using a large number of samples in every iteration. The second algorithm (Section 4.2) builds upon the first one by incorporating a technique from Blanchet and Glynn (2015) to reduce the bias of gradients evaluated at the private estimators of the iterates.
@@
  • In our algorithms, the iterates are privately estimated by sampling vertices proportionally to their weights, a technique known as Maurey’s Sparsification Lemma (Pisier, 1980-1981) [^3] This method is also known as the Maurey Empirical Method.. For nonlinear objectives, evaluating the duality gap and gradients of the objective at the iterate estimators introduces bias. To control the bias and other estimation errors, we provide various extensions of this classical result that work in terms of functional approximations. We state these results in more generality than what is needed for our analyses, as they could be of independent interest. See Section 3 for more details.
@@
  • We give the first mirror descent style algorithm for DP-SCO in the polyhedral setting with nearly dimension-independent excess risk guarantees by combining the ideas we used for DP-SSP with the Anytime Online-to-Batch Converstion from Cutkosky (2019). See Section 5. While these rates are not optimal, we expect they could inspire future work into more practical algorithms for this setting.
Setting
Objective
Previous best-known rate
Our rate
SSP
Bilinear
@@
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$ (Hardt and Rothblum, 2010)
@@
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$
(Thm. 4.1)
Quadratic
      None
@@
$\color[rgb]{1,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$
(Thm. 4.1)
$L_{2}$-smooth
      None
$\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{2/5}$
(Thm. 4.1)
$L_{1}$-smooth
      None
$\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/3}$
(Thm. 4.2)
$L_{2}$-smooth
      None
$\color[rgb]{1,0,0}\frac{\ell}{\sqrt{n}}+\frac{\ell}{\sqrt{n\varepsilon}}{}^{(\ast)}$
(Thm. 4.3)
SCO
$L_{2}$-smooth
      None
$\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}}{\sqrt{n\varepsilon}}$
(Thm. 5.1)
$L_{1}$-smooth
$\color[rgb]{1,0,0}\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{2/3}}{(n\varepsilon)^{2/3}}$ (Asi et al., 2021b)
$\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{7/10}}{(n\varepsilon)^{2/5}}$
(Thm. 5.1)
@@
Table 1: We omit polylog($n,1/\delta$) factors, function-class parameters and constants. We refer to Lipschitz and smooth functions as $L_{1}$-smooth. If the function has, in addition, smooth partial derivatives, then we refer to it as $L_{2}$-smooth. We denote $\ell=\log(d_{x})+\log(d_{y})$ and $\ell_{x}=\log(d_{x})$. Rates in red are optimal up to polylog factors. ${}^{(\ast)}$ This rate holds with constant probability while the rest hold in expectation.
@@ 1.1 Related Work
DP stochastic saddle-points.
@@
Stochastic (non-private) saddle-point problems have connections to optimization, economics, machine learning, and many other fields. One such algorithm for solving these problems is the stochastic entropic mirror-descent algorithm (Nemirovski et al., 2009). Recently, a line of research has studied DP stochastic saddle-points (Boob and Guzmán, 2023; Zhang et al., 2022a; Yang et al., 2022; Kang et al., 2022; Bassily et al., 2023). These works focus exclusively on Euclidean setups, which necessarily leads to a polynomial dependence on the dimension; furthermore, techniques of these works are entirely unrelated to ours. In addition, with only limited exceptions (Boob and Guzmán, 2023; Bassily et al., 2023), most works provide guarantees on the weak duality gap. Bassily et al. (2023) justify that the (strong) duality gap is a more meaningful and challenging efficiency estimate. This is the convergence criterion that we use in this work.
Optimization in the polyhedral setting.
@@
Closest to this work is DP stochastic convex optimization (SCO) in $\ell_{1}$ setup (Asi et al., 2021b; Bassily et al., 2021b; Han et al., 2022), as well as earlier work on private empirical risk minimization (ERM) (Jain and Thakurta, 2014; Talwar et al., 2014, 2015). These works attain nearly dimension-independent excess risk rates with private variants of the (stochastic) Frank-Wolfe method. Even though this method leads to optimal rates for DP-SCO, it is not known to yield any nontrivial rates for saddle-point problems. Even without DP constraints, it is not known how to solve SSP problems with Frank-Wolfe methods unless strong additional assumptions are made: namely, either simultaneously assuming smoothness and strong convexity/strong concavity [^4] These assumptions are indeed strong: If $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ is an $L_{1}$-smooth and $\mu$-strongly convex function w.r.t. $\|\cdot\|_{1}$, then $L/\mu=\Omega(d)$. This follows by conjugate duality and the argument in (d’Aspremont et al., 2018, Example 4.1), or strong convexity of the feasible sets (Gidel et al., 2017). For these reasons, those assumptions are of no use for the interests of this work.
Maurey Sparsification.
@@
This is a classical technique in high-dimensional convex geometry (Pisier, 1980-1981). The idea is based on sampling vertices with the barycentric coordinates weights and taking the empirical average, as nicely illustrated in the introduction of the book (Vershynin, 2018). This approximation is established for norms with nontrivial concentration estimates for independent random sums, e.g. quantified by their Rademacher type (Pisier, 2016). This technique has been used systematically as a compression-type argument to obtain faster rates in statistical estimation and learning (e.g. (Barron, 1993; Zhang, 2002)). More recently, Acharya et al. (2019) prove a result similar to Lemma 3.1, a functional approximation for randomized sparsification, in the context of model compression for distributed learning of generalized linear models. They also proposed a sparsified mirror descent algorithm that works in the $\ell_{1}/\ell_{q}$ setting where $q<\infty$, resorting on smoothnes properties of the Bregman divergence. Furthemore, they only sparsify the output models of each machine. We do not follow this approach. Instead, we sparsify every iterate of mirror descent. Furthermore, we show that entropic SMD (using a nonsmooth Bregman divergence) with iterate sparsification works for convex and smooth losses, which to our knowledge is an entirely new result, and does not follow from direct ex-post model compression.
Debiasing.
@@
This is a technique proposed in the simulation literature (Blanchet and Glynn, 2015), to leverage biased estimators whose accuracy improves with larger sample size. The main difference between the works on simulation and our approach is the finite sample size, which limits the amount of bias-reduction. Asi et al. (2021a) studied bias-reduction in regularization for SCO, but their results are unrelated to ours. In concurrent work (Ghazi et al., 2024), a gradient debiasing procedure is developed for Euclidean DP-SCO. Although our results do not depend on their techniques, we borrow some inspiration from them. Note futher that most of our results do not use bias-reduction.
Minimizing the worst-group loss.
@@
In independent and concurrent work, Zhou and Bassily (2024) investigate empirical and population worst-group risk minimization under DP, with a focus on group fairness and robust learning. While the convex-linear loss functions of their work correspond to a particular case of ours (convex-concave), their use of Euclidean geometry makes their results incomparable to ours. Furthermore, their sequential regularization approach is largely independent from ours.
2 Preliminaries

### v2

@@ Abstract
We study the problem of differentially-private (DP) stochastic (convex-concave) saddle-points in the $\ell_{1}$ setting. We propose $(\varepsilon,\delta)$-DP algorithms based on stochastic mirror descent that attain nearly dimension-independent convergence rates for the expected duality gap, a type of guarantee that was known before only for bilinear objectives. For convex-concave and first-order-smooth stochastic objectives, our algorithms attain a rate of $\sqrt{\log(d)/n}+(\log(d)^{3/2}/[n\varepsilon])^{1/3}$, where $d$ is the dimension of the problem and $n$ the dataset size. Under an additional second-order-smoothness assumption, we show that the duality gap is bounded by $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$ with high probability, by using bias-reduced gradient estimators. This rate provides evidence of the near-optimality of our approach, since a lower bound of $\sqrt{\log(d)/n}+\log(d)^{3/4}/\sqrt{n\varepsilon}$ exists. Finally, we show that combining our methods with acceleration techniques from online learning leads to the first algorithm for DP Stochastic Convex Optimization in the $\ell_{1}$ setting that is not based on Frank-Wolfe methods. For convex and first-order-smooth stochastic objectives, our algorithms attain an excess risk of $\sqrt{\log(d)/n}+\log(d)^{7/10}/[n\varepsilon]^{2/5}$, and when additionally assuming second-order-smoothness, we improve the rate to $\sqrt{\log(d)/n}+\log(d)/\sqrt{n\varepsilon}$. Instrumental to all of these results are various extensions of the classical Maurey Sparsification Lemma [Pisier:1980], which may be of independent interest.
keywords:
Differential Privacy, Stochastic Saddle Point Problem, Mirror Descent, Sparse Approximation
{MSCcodes}
90C47, 90C15, 62D99, 68P27
@@ 1 Introduction
In this work, we study the stochastic saddle-point (SSP) problem, that is, we aim to find a pair $(x^{*},y^{*})$ that solves
(1) $\min_{x\in\mathcal{X}}\max_{y\in\mathcal{Y}}\{F_{\mathcal{D}}(x,y):=\mathbb{E}_{z\sim\mathcal{D}}[f(x,y;z)]\},$
where the distribution $\mathcal{D}$ over the sample space $\mathcal{Z}$ is unknown, and only iid data from the target distribution, $S\sim\mathcal{D}^{n}$, are available. We assume that, for each $z\in\mathcal{Z}$, the function $(x,y)\to f(x,y;z)$ is convex-concave, and the function is $L_{0}$-Lipschitz and $L_{1}$-smooth w.r.t $\|\cdot\|_{1}$ (see Section 1.3 for details). We also assume that the feasible sets $\mathcal{X}$ and $\mathcal{Y}$ are the standard simplices in $\mathbb{R}^{d_{x}}$ and $\mathbb{R}^{d_{y}}$, denoted by $\Delta_{x}$ and $\Delta_{y}$ from now on [^1] Note that we can reduce SSP problems over polytope feasible sets to this setting. In this case, the dimensions $d_{x}$ and $d_{y}$ will correspond to the number of vertices of $\mathcal{X}$ and $\mathcal{Y}$, respectively.. We will refer to SSP problems satisfying these assumptions as the $\ell_{1}$ setting’ [JNLS:2009, NesterovNemirovski:2013]. See Section 1.4 for important problems where these assumptions hold.
In modern machine learning applications, the observed dataset $S$ usually contains sensitive user information. In this context, Differential Privacy (DP) has become the de facto standard for privacy-preserving algorithms. A randomized algorithm $\mathcal{A}(S)$ is $(\varepsilon,\delta)$-differentially private if for any pair of datasets $S$ and $S^{\prime}$ differing in at most one data point and any event $\mathcal{E}$ in the output space,
$\mathbb{P}[\mathcal{A}(S)\in\mathcal{E}]\leq e^{\varepsilon}\mathbb{P}[\mathcal{A}(S^{\prime})\in\mathcal{E}]+\delta.$
We study the Differentially Private Stochastic Saddle Point (DP-SSP) problem in the $\ell_{1}$ setting. More concretely, we provide differentially private algorithms that receive $S$ as input and return a pair $(x_{\mathcal{A}(S)},y_{\mathcal{A}(S)})$ that approximates an SSP. Given a pair $(x,y)$, we measure its efficiency by the duality gap, defined as
(2) $\operatorname{Gap}(x,y)=\max_{v\in\mathcal{X},w\in\mathcal{Y}}(F_{\mathcal{D}}(x,w)-F_{\mathcal{D}}(v,y)).$
There is a major discrepancy in the literature between the achievable rates of convergence for the duality gap between some structured problems and more general convex-concave settings . While some bilinear problems have convergences rates which scale only poly-logarithmically with the ambient dimensions
@@
(e.g.,  [Hsu:2013]), all known rates for general convex-concave objectives scale polynomially with the dimension (e.g., [pmlr-v195-bassily23a]). This can be attributed to the fact that all existing results for convex-concave settings work under Euclidean geometry, which assumes general convex and compact feasible sets and objectives that are Lipschitz and smooth with respect to $\|\cdot\|_{2}$, as opposed to the $\ell_{1}$ setting considered in this work. Therefore, known polynomial dimension lower bounds for DP Stochastic Convex Optimization (SCO) in the Euclidean setting [Bassily:2014] apply. Recently, [asi2021private, bassily2021differentially] showed that it is possible to obtain excess risk bounds for DP-SCO under $\ell_{1}$ setting which are nearly independent of the dimension. Inspired by these recent developments, in this work we provide DP algorithms with nearly dimension-independent rates for convex-concave SSP problems in the $\ell_{1}$ setting. Our results extend considerably the class of SSP problems where it is possible to bypass polynomial in the dimension lower bounds, which so far only contain problems with bilinear objectives and simplex feasible sets.
1.1 Summary of Contributions
We prove the first nearly-dimension independent rates for convex-concave DP-SSP in the $\ell_{1}$ setting. These rates are optimal (up to logarithmic factors) for second-order-smooth objectives, and have the form $r_{NDP}+r_{DP}$ where $r_{NDP}$ is the optimal non-private SSP rate and $r_{DP}$ is an error rate due to differential privacy. Furthemore, the algorithms achieving these rates have an $O(n)$ gradient complexity. We also use our techniques to show nearly-dimension independent convergence rates for DP-SCO under the same assumptions considered in [asi2021private, bassily2021differentially]. Previous work used private versions of the variance-reduced stochastic Frank-Wolfe algorithm. By contrast, we are the first to achieve this using variants of mirror-descent algorithms. Table 1 summarizes the obtained rates.
@@ 1.1 Summary of Contributions
  • We propose two algorithms for DP-SSP, based on (entropic) Stochastic Mirror Descent (SMD). As opposed to previous results for bilinear settings where privatization by sampling from vertices leads to unbiased gradient estimators [Hsu:2013], the main challenge we encounter is controlling the bias in the more general (stochastic) convex-concave setting. The first algorithm (Section 3.1) controls bias by simply using a large number of samples in every iteration. The second algorithm (Section 3.2) builds upon the first one by incorporating a stochastic simulation technique from [Blanchet:2015] to reduce the bias of gradients evaluated at the private estimators of the iterates.
@@
  • In our algorithms, the iterates are privately estimated by sampling vertices proportionally to their weights, a technique known as Maurey’s Sparsification Lemma [Pisier:1980] [^2] This method is also known as the Maurey Empirical Method.. We provide various extensions of this classsical result, most notably for the duality gap and gradient sampling from vertices, that are new and we believe are of independent interest. These extensions are challenging, primarily due to the presence of bias in the underlying estimation task. See Section 2 and Subsection 3.1.2 for more details.
@@
  • We give the first mirror descent style algorithm for DP-SCO in the polyhedral setting with nearly dimension-independent excess risk guarantees by combining the ideas we used for DP-SSP with the Anytime Online-to-Batch Conversion from [anytime-online-to-batch]. See Section 4. While these rates are not optimal, we expect they could inspire future work into more practical algorithms.
Setting Objective Previous best-known rate Our rate
@@
SSP Bilinear $\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}\text{\tiny{(*)}}$ [Hardt:2010] $\color[rgb]{0,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$(*) (Thm. 3.1)
@@
Quadratic     None $\color[rgb]{0,0,0}\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/2}$(*) (Thm. 3.1)
@@
$L_{2}$-smooth     None $\color[rgb]{0,0,0}\frac{\ell}{\sqrt{n}}+\left(\frac{\ell^{2}}{n\varepsilon}\right)^{1/2}$(**) (Thm. 3.14)
$L_{1}$-smooth     None $\frac{\sqrt{\ell}}{\sqrt{n}}+\left(\frac{\ell^{3/2}}{n\varepsilon}\right)^{1/3}$ (Thm. 3.1)
SCO $L_{2}$-smooth     None $\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}}{\sqrt{n\varepsilon}}$ (Thm. 4.2)
$L_{1}$-smooth $\color[rgb]{0,0,0}\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{2/3}}{(n\varepsilon)^{2/3}}$(*) [asi2021private] $\frac{\sqrt{\ell_{x}}}{\sqrt{n}}+\frac{\ell_{x}^{7/10}}{(n\varepsilon)^{2/5}}$ (Thm. 4.2)
@@
Table 1: Rates derived from our results. We omit polylog($n,1/\delta$) factors, function-class parameters and constants. We refer to first-order-smooth and second-order-smooth functions as $L_{1}$-smooth and $L_{2}$-smooth, respectively. We denote $\ell=\log(d_{x})+\log(d_{y})$ and $\ell_{x}=\log(d_{x})$. (*): these rates are optimal up to polylog factors. (**): this rate holds with high probability while the rest hold in expectation.
@@ 1.2 Related Work
DP stochastic saddle-points
@@
Stochastic (non-private) saddle-point problems have connections to optimization, economics, machine learning, and many other fields. One such algorithm for solving these problems is the stochastic entropic mirror-descent algorithm [JNLS:2009], which is the basis for much of this work. Recently, a line of research has studied the differentially private (DP) version of stochastic saddle-point problems [boob2024optimal, DPSMO, DPSGDA, Kang:2022, pmlr-v195-bassily23a]. These works focus exclusively on Euclidean setups, which necessarily leads to algorithms that have a polynomial dependence on the dimension. Moreover, due to the Euclidean setup, the techniques used in these works do not necessarily apply to the setting considered here. In addition, with only limited exceptions [boob2024optimal, pmlr-v195-bassily23a], most previous work provide guarantees on the weak duality gap. The (strong) duality gap (2), on the other hand, is more meaningful (see [pmlr-v195-bassily23a] for justification), but also more difficult to bound, most notably in the private setting. This is the convergence criterion that we use in this work.
Optimization in the $\ell_{1}$ setting
@@
Closest to this work is DP stochastic convex optimization (SCO) in the $\ell_{1}$ setting [asi2021private, bassily2021non, Han:2022], as well as earlier work on private empirical risk minimization (ERM) [Jain:2014, Talwar:2014, Talwar:2015]. These works attain nearly dimension-independent excess risk rates with private variants of the (stochastic) Frank-Wolfe method. Even though this method leads to optimal rates for DP-SCO, it is not known to yield any nontrivial rates for saddle-point problems. Even without DP constraints, it is not known how to solve SSP problems with Frank-Wolfe methods unless strong additional assumptions are made: namely, either simultaneously assuming smoothness and strong convexity/strong concavity [^3] These assumptions are indeed strong: If $f:\mathbb{R}^{d}\mapsto\mathbb{R}$ is an $L_{1}$-smooth and $\mu$-strongly convex function w.r.t. $\|\cdot\|_{1}$, then $L_{1}/\mu=\Omega(d)$. This follows by conjugate duality and the argument in [dAspremont:2018, Example 4.1], or strong convexity of the feasible sets [Gidel:2017]. Clearly, these assumptions are not suitable for this work.
Maurey Sparsification
@@
This is a classical technique in high-dimensional convex geometry [Pisier:1980]. The idea is based on sampling vertices with the barycentric coordinates weights and taking the empirical average, as nicely illustrated in the introduction of the book [Vershynin:2018]. This approximation is established for norms with concentration estimates for independent random sums, e.g., quantified by their Rademacher type [Pisier:2016]. This technique has been used systematically as a compression-type argument to obtain faster rates in statistical estimation and learning (e.g.,  [Barron:1993, Zhang:2002]). More recently, [Acharya:2019] proved a result similar to Lemma 2.1, a functional approximation for randomized sparsification, in the context of model compression for distributed learning of generalized linear models. They also proposed a sparsified mirror descent algorithm that works in the $\ell_{1}/\ell_{q}$ setting where $q<\infty$, resorting on smoothness properties of the Bregman divergence. Furthemore, they only sparsify the outputs of each machine. We do not follow this approach. Instead, we sparsify every iterate of mirror descent. Furthermore, we show that entropic SMD (using a nonsmooth Bregman divergence) with iterate sparsification works for convex and smooth losses, which to our knowledge is an entirely new result, and does not follow from direct ex-post model compression.
Debiasing
@@
This is a technique proposed in the stochastic simulation literature [Blanchet:2015], to leverage biased estimators whose accuracy improves with larger sample size. The main difference between the works on simulation and our approach is the finite sample size, which limits the amount of bias-reduction. The work [Asi:2021bias] studied bias-reduction in regularization for SCO. In concurrent work [Ghazi:2024], a gradient debiasing procedure is developed for Euclidean DP-SCO. Although our results do not depend on their techniques, we borrow some inspiration from them. Note futher that most of our results do not use bias-reduction.
Concurrent and follow up work
@@
In independent and concurrent work, the paper [Zhou:2024] investigates empirical and population worst-group risk minimization under DP, with a focus on group fairness and robust learning. While the convex-linear loss functions of their work correspond to a particular case of ours (convex-concave), their use of Euclidean geometry makes their results incomparable to ours. Furthermore, their sequential regularization approach is independent from ours. The follow up work [BGM24noneuclideanDP-SSP] considers DP-SSPs in $\ell_{p}/\ell_{q}$ setup with $p,q\in[1,2]$. In contrast to our work, their results focus exclusively in the nonsmooth setting, leading to unavoidable polynomial-in-the-dimension convergence rates for the strong duality gap.
1.3 Preliminaries